Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-05-06

Observers: Sandra Benitez, Joe Polshaw
Support: Cosimo Inserra
TAT on duty: Maria Teresa Botticella, Emma Walker

--------------------------------------------------------------

WEATHER CONDITIONS:
start: wind 0.6 m/s; temp. 14.1 C; humid. 13.0%; Mostly clear, some cirrus
middle: wind 5.0 m/s; temp. 14.9 C: humid. 12.0%; Mostly clear
end: wind 4.7 m/s; temp. 15; humid. 11%; Thin clouds


--------------------------------------------------------------

INSTRUMENTS: EFOSC2, SOFI

CALIBRATIONS

Afternoon:
-- EFOSC2 bias, dome flats, arcs

Morning:
-- EFOSC arcs

--------------------------------------------------------------

OBSERVATIONS

Target                   Type                         UT-Start         Filter/Grism           Slit (")         Seeing (")      Notes

SN2013ai              Follow (IMAG)            23:18               VRi                        N/A           0.6 - 0.9       (1)
LSQ14bdq             Follow (SPEC)             00:25              Gr11                      1               0.6 - 1.4       (2)
CSS140424...        Follow (SPEC)             01:17              Gr11                      1               0.9 - 1.2       (3)
CSS140424...        Follow (IMAG)            02:30              UBVRi                     N/A               0.7            (4)
LTT3864               STD (SPEC)                02:46               Gr11/13/16          1/1.5         0.7                -
CSS140421...        Follow (SPEC)            03:16               Gr13                       1              0.9 - 1.2        (5)

*** CHANGE TO SOFI ***

CSS140421...        Follow (IMAG)            04:49               JHKs                       N/A              1.2             (6)
CSS140424...        Follow (IMAG)            05:56               JHKs                       N/A              1.0 - 1.2    (7)

*** CHANGE TO EFOSC2 ***

SN2009ip             Follow (SPEC)            07:08                 Gr16                       1             1.1 - 1.4       (8)
OGLE-79              Follow (SPEC)            08:56                 Gr13                       1             1.1 - 1.3        (9)
SN2013fs             Follow (IMAG)           10:07                 VRi (i useless!)        N/A         1.0 - 1.5       (10)


(1) 6*150s VRi.
(2) We have reduced the exposure time from 2*1800s to 1*2400s here. This is because we have a *very* tight schedule tonight to fit in as many of the followup requests as possible for the night, and the priority for this target was listed as USEFUL (whereas the later targets are either CRITICAL or IMPORTANT). A bit of time lost since the telescope had to be focused. There were large fluctuations in the seeing during the OB.
(3) CSS140424-133007-212728. 2*1800s gr11
(4) CSS140424-133007-212728. 300s U, 45s B, 30s V, 20s Ri
(5) CSS140421-142042+031602. 2*1800s gr13. Exposure time reduced from 3*1800s to allow more time for SOFI imaging. Will do EFOSC UBVRi after SOFI if we have time, but there are many high priority targets at the end of the night.
(6) CSS140421-142042+031602. Lost quite a lot of time here since there was a problem with the image analysis. The exposure times have been slightly reduced from the previous OB since we are behind schedule... 900s J, 1080s H, 1080s Ks.
(7) CSS140424-133007-212728. 720s JH, 1080s Ks. The exposure times have been increased from the previous OB (4th May) but not as long as we would have liked due to the tight schedule.
(8) Began at high airmass. 3*1800s Gr16 - reduced from the requested 4*1800s in order to fit OGLE-079 and SN2013fs in (both also CRITICAL priority targets). Unfortunately there is no time to do the imaging of CSS140421-142042+031602 or SSS120810-231802-560926 (the latter was given IMPORTANT priority). We also would have liked to observe another STD here, but since we already observed one earlier in the night we favoured ensuring we got as close to fulfilling the followup requests as possible.
(9) 3600s Gr13 spectroscopy of the nearby red source.
(10) 3*240s V, 3*120s Ri. Observing into the start of twilight. Too late to do another STD... Unfortunately the i-band is completely saturated!!! The V and R bands should be fine though.

Comments